------------------------------------------------------------
Date: Sat Oct 27 08:18:43 2001, Shift: 00:00-08:00
Supervisor: JC, with: 
Log book#10 pages 61 - 62
Shift Summary:
------------------------------------------------------------
1. 
Quite a good store continued from the previous shift. 
Around 2 am Brahms and Phenix voted to dump. 
Rhic delivered a new beam at 4:45 but 15 minutes later 
the beam was dumped abruptly without warning (run 5534). 
Reason: "Cryogenic problems. Cryo permit pulled by 
leadflow problem".
2. 
Three data runs:
              Trig 1     Trig 6
  Run 5528    192k         60k
  Run 5529     93k	   25k
  Run 5534     70k	   22k  	
Pedestals and DVM runs: 5531-5533.
3.
Dead channel in TOFW Top tube 92.
No other detector problems.
------------------------------------------------------------
This archive was generated by hypermail 2b30 : Sat Oct 27 2001 - 08:19:14 EDT